Solve the quadratic equation using the method that seems most appropriate to you.
Also, how do you know which method is best?
n^2 + 16n +63 = 0
Thanks in advance!

Answer by ewatrrr(24785)


Hi,
Solving the quadratic equation by factoring
n^2 + 16n +63 = 0
(n+7)(n + 9) = 0
(n+7)= 0
n = -7
(n + 9) = 0
n = -9